OpenStudy (anonymous):

solve for x , 2x^2x = 9^x-3

OpenStudy (dumbcow):

can you clarify exponents with parenthesis?

OpenStudy (anonymous):

2x^(2x) = 9^(x-3)

OpenStudy (dumbcow):

hmm since you have a x^x term , im not sure if this solvable using algebra

OpenStudy (dumbcow):

take log both sides \[2x \ln x = (x-3) \ln 9 - \ln 2\] \[x(2\ln x - \ln9) = -3\ln 9 -\ln2 \] ...no way to isolate "x" since it both inside log and outside

OpenStudy (anonymous):

ok thank you

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=2x%5E%282x%29+%3D+9%5E%28x-3%29 its ugly...may be no solution